# Continued Fractions and the $4$-Color Theorem

### v6

@@ Abstract
We study the geometry of some proper $4$-colorings of the vertices of sphere triangulations with degree sequence $6,...,6,2,2,2$. Such triangulations are the simplest examples which have non-negative combinatorial curvature. The examples we construct, which are roughly extremal in some sense, are based on a novel geometric interpretation of continued fractions. We will also present a conjectural sharp “isoperimetric inequality” for colorings of this kind of triangulation.
@@ 1.1 Background
A triangulation of non-negative combinatorial curvature is one in which the maximum degree is $6$. All the vertices have degree $6$ except for a list $v_{1},...,v_{k}$ which have degrees $62\leq d_{1},...,d_{k}<$ 6. Euler’s Formula gives the condition on the degrees:
@@ 1.2 The Continued Fraction Colorings
The $6,...,6,2,2,2$ trianglulations are indexed by the nonzero Eisenstein integers. An Eisenstein integer is a number of the form
$a+b\alpha,\hskip 30.0pta,b\in\mbox{\boldmath{$Z$}},\hskip 15.0pt\alpha=\frac{1+i\sqrt{3}}{2}.$
(It is more common to use $\omega=\alpha^{2}$ in place of $\alpha$, but $\alpha$ is more convenient for us.) To see the connection, let $\cal E$ denote the ring of Eisenstein integers. The points of $\cal E$ are naturally the vertices of an equilateral triangulation $\cal T$ of $C$. Given some nonzero $\beta\in\cal E$, we let $\beta\cal T$ be the bigger equilateral triangulation obtained by multiplying the whole picture by $\beta$. We let $G_{\beta}$ denote the group of symmetries generated by order $3$ rotations in the vertices of $\beta\cal T$. The quotient
@@ 1.3 Properties of the Continued Fraction Colorings
Here $\phi$ is the golden ratio. Our proof of Statement 2 of Theorem 1.1 will show more generally that $\eta(\zeta)\in\mbox{\boldmath{$Q$}}(\zeta)$, the quadratic field containing $\zeta$. See §3.3 for some examples.
@@
Let $\cal C$ be any good coloring of any triangulation with degree sequence $6,...,6,2,2,2$. Then $\eta({\cal C\/})\geq 3$.
@@
Suppose $\{{\cal G\/}_{n}\}$ is any infinite sequence of distinct colorings of primitive sphere triangulations with degree sequence $6,...,6,2,2,2$. Then $\liminf_{n\to\infty}\eta({\cal G\/}_{n})\geq\phi^{6}$.
@@
There is a constant $\Omega$ with the following property. For any triangulation $\cal T$ with degree sequence 6,.,6,2,2,26,....,6,2,2,26 , … . , 6 , 2 , 2 , 2 there is a good coloring $\cal C$ of $\cal T$ with $\eta({\cal C\/})<\Omega$.
@@ 2.2 Trapezoid Necklaces
An isosceles trapezoid is a quadrilateral with two parallel sides such that the other two sides are non-parallel but have the same length. We call the longer parallel side the top, the shorter parallel side the bottom, and the other two sides the diagonal sides. We allow the degenerate case of an isosceles triangle. In this case the bottom has length 00. An Eisenstein trapezoid is an isosceles trapezoid whose edges lie in the $1$-skeleton of $\cal T$, the planar equilateral triangulation whose vertices are the Eisenstein integers. Figure 1 above and Figure 3 below feature some Eisenstein trapezoids.
Up to symmetries of $\cal T$, an Eisenstein trapezoid $X$ is characterized by the pair $(a,b)$ where $a$ is the length of a diagonal side of $X$ and $b$ is the length of the top. We call $X$ primitive if $a,b$ are relatively prime. When $X$ is primitive, we define the aspect ratio to be $a/b$. The aspect ratio determines the primitive Eisenstein trapezoid up to symmetries of $\cal T$. Thus, modulo symmetry, the primitive Eisenstein trapezoids are naturally in bijection with the set $\cal S$ of rationals considered above. The Eisenstein trapezoids in Figure 1 are all primitive, and their aspect ratios are variously $/1$ 1 and $1/2$ and $2/3$. The Eisenstein trapezoids in Figure 3 have aspect ratio $3/5$.
@@
In this description the indices are taken mod $6$. We define the center of the necklace to be the fixed point of $\rho$. When the center is 00, the map $\rho$ (or perhaps its inverse) is multiplication by $\alpha$. We define the aspect ratio of the necklace to be the common aspect ratio of the $6$ individual trapezoids.
Up to symmetry of $\cal T$, there exists a unique Eisenstein necklace of aspect ratio $a/b\in\cal S$. If we normalize the picture so that 00 is the center, then one of the trapezoids $X_{1}$ has vertices
[…]
@@ 2.3 Empty Trapezoid Flowers
In general, the empty $p/q$-flower starts with a $p/q$-necklace and then fills in the full $\gamma$ orbit, alternately coloring the necklaces black and white. The innermost necklace always has aspect ratio $/1$ 1. Up to symmetries of $\cal T$, the empty $p/q$-flower is unique. One can read off the continued fraction expansion of $p/q$ by counting the stripes of the maximal trapezoids.
@@ 3.1 Zero Degree
Let $T$ denote the regular tetrahedron. Call a triangulation of the sphere even if it has all even degrees. In this section we prove that any good coloring of an even triangulation $\Sigma$ has the same total number of black and white triangles. This is equivalent to the statement that the associated map $f:\Sigma\to T$ has topological degree 00. I learned this proof from Kasra Rafi.
[…]
@@ 3.2 Asymptotic Properties
Statement 1: Let $G(p_{n}/q_{n})$ denote the sequence of rationals we get by starting with $p_{n}/q_{n}$, iteratively applying the slow Gauss map, then listing the terms in reverse order. For instance,
[…]
@@
For any $D$ there are constants $M,N$ such that if $n>N$ then all but the first $M$ terms of $G(p_{n}/q_{n})$ have $p_{n}>D$.
Proof: Our sequences stabilize in the following sense. For any given $M$, we can take $N$ sufficiently large so that the first $M$ terms of $G(p_{n}/q_{n})$ are the same as long as $n>N$. What makes it work is that the thinner sequence of continued fraction approximants of $p_{n}/q_{n}$ converges to the sequence of continued fraction approximants of $\zeta$, and we can derive the $G$-sequences from these thinner sequences.
We also note a second property of our sequences: The numerators of the terms in $G(p_{n}/q_{n})$ are monotone increasing as we move along the sequence. When we combine the increasing numerator property with the stabilization property, we get the result of this lemma. $\spadesuit$
In terms of the flowers, Lemma 3.1 says that all but the first $M$ inner trapezoid necklaces $Q$ have the property that
$\frac{p(Q)}{A(Q)}<\frac{100}{D}.$
Here $p(Q)$ denotes the perimeter of $Q$ and $A(Q)$ denotes the number of equilateral triangles comprising $Q$. We picked an unrealistically large constant of $100$ here to avoid having to think about the fine points of trapezoids.
Our analysis shows that, within the $n$th flower, the average ratio of the perimeter of a trapezoid to the area of the trapezoid tends to 00 as $n\to\infty$. This immediately implies that the ratio $f_{n}/F_{n}$ converges to 00.
@@
[…]
@@
[…]
@@ 4.1 Special Hexagons
Say that a special hexagon is a convex hexagon whose interior angles are all $32\pi/$ 3. The regular hexagon is an example, and all other special hexagons are obtained from the regular one by pushing the sides in and out parallel to themselves. Our first result is a version of the isoperimetric inequality for special polygons. This result is undoubtedly well-known.
@@ 4.2 The Main Argument
Let $\cal C$ be a good coloring of some triangulation $\cal T$ with degree sequence $6,...,6,2,2,2$. Let $f$ denote the fold count and let $F$ be the number of faces. Our goal is to show that $\eta({\cal C\/})=f^{2}/F\geq 3$. Note that $\cal C$ partitions $\cal T$ into a finite number of monochrome regions with connected interior.
@@
Consider a lift $\widetilde{P}_{\epsilon}$ of $P_{\epsilon}$ to $C$, the orbifold universal cover of our flat cone sphere. This lift is also locally convex near its boundary components. If $\widetilde{P}_{\epsilon}$ is unbounded then the local convexity near the boundary forces it to be an infinite strip. But then $P_{\epsilon}$ is an annulus with geodesic boundary. This is impossible in our flat cone sphere. Hence $\widetilde{P}_{\epsilon}$ is bounded. The local convexity forces $\widetilde{P}_{\epsilon}$ to have just one boundary component. If the map $\widetilde{P}_{\epsilon}\to P_{\epsilon}$ is not a global isometry then $P_{\epsilon}$ contains some degree $2$ vertices, a contradiction. We conclude that $\widetilde{P}_{\epsilon}$ is a convex polygon and the map $\widetilde{P}_{\epsilon}\to P_{\epsilon}$ is an isometry. Hence $P_{\epsilon}$ is a convex polygon.

### v7

@@ Abstract
We study the geometry of some proper $4$-colorings of the vertices of sphere triangulations with degree sequence 6,,6,2,2,26,...,6,2,2,2. Such triangulations are the simplest examples which have non-negative combinatorial curvature. The examples we construct, which are roughly extremal in some sense, are based on a novel geometric interpretation of continued fractions. We will also present a conjectural sharp “isoperimetric inequality” for colorings of this kind of triangulation.
@@ 1.1 Background
A triangulation of non-negative combinatorial curvature is one in which the maximum degree is $6$. All the vertices have degree $6$ except for a list $v_{1},...,v_{k}$ which have degrees $2\leq d_{1},...,d_{k}<6$. Euler’s Formula gives the condition on the degrees:
@@ 1.2 The Continued Fraction Colorings
The 6,,6,2,2,26,...,6,2,2,2 trianglulations are indexed by the nonzero Eisenstein integers. An Eisenstein integer is a number of the form
$a+b\alpha,\hskip 30.0pta,b\in\mbox{{$Z$}},\qquad\alpha=\frac{1+i\sqrt{3}}{2}.$
[…]
@@ 1.3 Properties of the Continued Fraction Colorings
Here $\phi$ is the golden ratio. Our proof of Statement 2 of Theorem 1.1 will show more generally that $\eta(\zeta)\in\mbox{{$Q$}}(\zeta)$, the quadratic field containing $\zeta$. See §3.3 for some examples.
@@
Let $\cal C$ be any good coloring of any triangulation with degree sequence 6,,6,2,2,26,...,6,2,2,2. Then $\eta({\cal C\/})\geq 3$.
@@
Suppose $\{{\cal G\/}_{n}\}$ is any infinite sequence of distinct colorings of primitive sphere triangulations with degree sequence 6,,6,2,2,26,...,6,2,2,2. Then $\liminf_{n\to\infty}\eta({\cal G\/}_{n})\geq\phi^{6}$.
@@
There is a constant $\Omega$ with the following property. For any triangulation $\cal T$ with degree sequence 6,.,6,2,2,26,....,6,2,2,2 there is a good coloring $\cal C$ of $\cal T$ with $\eta({\cal C\/})<\Omega$.
@@ 2.2 Trapezoid Necklaces
An isosceles trapezoid is a quadrilateral with two parallel sides such that the other two sides are non-parallel but have the same length. We call the longer parallel side the top, the shorter parallel side the bottom, and the other two sides the diagonal sides. We allow the degenerate case of an isosceles triangle. In this case the bottom has length 0. An Eisenstein trapezoid is an isosceles trapezoid whose edges lie in the $1$-skeleton of $\cal T$, the planar equilateral triangulation whose vertices are the Eisenstein integers. Figure 1 above and Figure 3 below feature some Eisenstein trapezoids.
Up to symmetries of $\cal T$, an Eisenstein trapezoid $X$ is characterized by the pair $(a,b)$ where $a$ is the length of a diagonal side of $X$ and $b$ is the length of the top. We call $X$ primitive if $a,b$ are relatively prime. When $X$ is primitive, we define the aspect ratio to be $a/b$. The aspect ratio determines the primitive Eisenstein trapezoid up to symmetries of $\cal T$. Thus, modulo symmetry, the primitive Eisenstein trapezoids are naturally in bijection with the set $\cal S$ of rationals considered above. The Eisenstein trapezoids in Figure 1 are all primitive, and their aspect ratios are variously $1/1$ and $1/2$ and $2/3$. The Eisenstein trapezoids in Figure 3 have aspect ratio $3/5$.
@@
In this description the indices are taken mod $6$. We define the center of the necklace to be the fixed point of $\rho$. When the center is 0, the map $\rho$ (or perhaps its inverse) is multiplication by $\alpha$. We define the aspect ratio of the necklace to be the common aspect ratio of the $6$ individual trapezoids.
Up to symmetry of $\cal T$, there exists a unique Eisenstein necklace of aspect ratio $a/b\in\cal S$. If we normalize the picture so that 0 is the center, then one of the trapezoids $X_{1}$ has vertices
$a+b\alpha,\hskip 30.0pt(a-b)+b\alpha,\hskip 30.0pt(2a-b)+(b-a)\alpha,\hskip 30.0pta+(b-a)\alpha.$
@@ 2.3 Empty Trapezoid Flowers
In general, the empty $p/q$-flower starts with a $p/q$-necklace and then fills in the full $\gamma$ orbit, alternately coloring the necklaces black and white. The innermost necklace always has aspect ratio $1/1$. Up to symmetries of $\cal T$, the empty $p/q$-flower is unique. One can read off the continued fraction expansion of $p/q$ by counting the stripes of the maximal trapezoids.
@@ 3.1 Zero Degree
Let $T$ denote the regular tetrahedron. Call a triangulation of the sphere even if it has all even degrees. In this section we prove that any good coloring of an even triangulation $\Sigma$ has the same total number of black and white triangles. This is equivalent to the statement that the associated map $f:\Sigma\to T$ has topological degree 0. I learned this proof from Kasra Rafi.
We equip $\mbox{{$R$}}^{2}$ with the unit equilateral triangulation. Let $\pi:\mbox{{$R$}}^{2}\to T$ be the branched covering which maps triangles to triangles in an isometric way. A good triangulation has the property that, around each vertex, the number of black triangles is congruent mod $3$ to the number of white triangles. In the case of an even triangulation, a good coloring has the stronger property that around each vertex the number of black triangles is congruent mod $6$ to the number of white triangles. But this means that the map $f:\Sigma\to T$ lifts to a map $\widetilde{f}:\Sigma\to\mbox{{$R$}}^{2}$ such that $f=\pi\circ\widetilde{f}$. The map $\widetilde{f}$ is homotopic to a point and hence so is $f$. But then $f$ has topological degree 0.
@@ 3.2 Asymptotic Properties
Statement 1: If we have an isosceles trapezoid based on the pair $(a,b)$ then the length-to-area ratio is at most $100/a$. This is conservative estimate. We fix $D$ for the whole discussion and show that $f_{n}/F_{n}<200/D$ once $n$ is large enough. Letting $D\to\infty$ we see that $f_{n}/F_{n}\to 0$.
Let $G(p_{n}/q_{n})$ denote the sequence of rationals we get by starting with $p_{n}/q_{n}$, iteratively applying the slow Gauss map, then listing the terms in reverse order. For instance,
$G(3/7)=\frac{1}{1},\quad\frac{1}{2},\quad\frac{1}{3},\quad\frac{3}{4},\quad\frac{3}{7}.$
We call this the $G$-sequence.
Let $p^{\prime}_{n}/q^{\prime}_{n}$ be the term furthest along in the $G$-sequence with $p_{n}^{\prime}\leq D$. Let $R_{n}$ be the radius of the flower based on $p_{n}/q_{n}$ and likewise define $R^{\prime}_{n}$.
@@
$R_{n}^{\prime}/R_{n}\to 0$ as $n\to\infty$.
Proof: There is a uniform constant $C>0$ such that
$\frac{1}{C}<\frac{R_{n}}{p_{n}+q_{n}},\frac{R_{n}^{\prime}}{p_{n}^{\prime}+q_{n}^{\prime}}<C.$ (7)
So, it suffices to prove that $\rho_{n}:=(p_{n}^{\prime}+q_{n}^{\prime})/(p_{n}+q_{n})\to 0$ as $n\to\infty$.
Say that a big step happens in the $G$-sequence whenever the numerator for consecutive terms changes. Whenever a big step happens, the number $a/b$ precedes the number $b/(a+b)$. The ratio of the sums is $(a+b)/(a+2b)<2/3$. So, if $m_{n}$ big steps are taken between $p^{\prime}_{n}/q^{\prime}_{n}$ and $p_{n}/q_{n}$ then $\rho<(2/3)^{m_{n}}$.
Let $M_{n}$ be the total number of big steps taken in $G(p_{n}/q_{n})$. Note that $m_{n}\geq M_{n}-D$. By continuity, the Gauss map for $p_{n}/q_{n}$ matches the Gauss map for the limit $\zeta$ for a number of initial steps that tends to $\infty$ as $n\to\infty$. Hence $M_{n}\to\infty$ as $n\to\infty$. Hence $m_{n}\to\infty$ as $n\to\infty$. Hence $\rho_{n}\to 0$. $\spadesuit$
Remark: I am grateful to Emma Rüter for pointing out a flaw in the version of Lemma 3.1 that I had in an earlier version of this paper.
The little flower based on $p_{n}^{\prime}/q_{n}^{\prime}$ is a subset of the big flower based on $p_{n}/q_{n}$. We have
$f_{n}^{\prime}<100F_{n}^{\prime},\hskip 30.0pt\frac{f_{n}-f_{n}^{\prime}}{F_{n}-F_{n}^{\prime}}\leq\frac{100}{D}.$ (8)
The first bound holds in any flower. The second bound comes from the fact that we are only considering trapezoids outside the small flower, and these all have length-to-area ratio at most $100/D$.
For any $\epsilon>0$ we can take $n$ so large that $R_{n}^{\prime}/R_{n}<\epsilon$. But then, by scaling, we have $F_{n}^{\prime}/F_{n}<\epsilon^{2}$. Choosing $n$ sufficiently large, so as to make the final inequality true, we have
$\frac{f_{n}}{F_{n}}=\frac{(f_{n}-f_{n}^{\prime})+f_{n}^{\prime}}{F_{n}}\leq\frac{(f_{n}-f_{n}^{\prime})+100F_{n}^{\prime}}{F_{n}}\leq$
$\frac{(f_{n}-f_{n}^{\prime})}{F_{n}}+100\epsilon^{2}<\frac{f_{n}-f^{\prime}_{n}}{F_{n}-F^{\prime}_{n}}+100\epsilon^{2}\leq\frac{100}{D}+100\epsilon^{2}<\frac{200}{D}.$ (9)
This proves Statement 1.
@@
We think of ${\cal T\/}(p_{n}+q_{n}\alpha)$ as a coloring of a doubled equilateral triangle. We then scale the metric by a factor of $q_{n}^{-1}$. The corresponding sequence of doubled equilateral triangles converges (say, in the Gromov-Hausdorff topology) to another doubled equilateral triangle, namely $\mbox{{$C$}}/G_{\beta}$ where $\beta=\zeta+\alpha.$ Again, $G_{\beta}$ is the group generated by order $3$ rotations in points of $\beta{\cal E\/}$.
@@
(1) The quantities $C_{1}\sqrt{3}$ and $C_{2}\sqrt{3}$ both belong to $\mbox{{$Q$}}(\zeta)$. Likewise $C_{3},C_{4},\lambda$ also all belong to $\mbox{{$Q$}}(\zeta)$. Therefore, the limiting value $\eta(\zeta)$ lies in $\mbox{{$Q$}}(\zeta)$. I will give some calculations below.
@@ 4.1 Special Hexagons
Say that a special hexagon is a convex hexagon whose interior angles are all $2\pi/3$. The regular hexagon is an example, and all other special hexagons are obtained from the regular one by pushing the sides in and out parallel to themselves. Our first result is a version of the isoperimetric inequality for special polygons. This result is undoubtedly well-known.
@@ 4.2 The Main Argument
Let $\cal C$ be a good coloring of some triangulation $\cal T$ with degree sequence 6,,6,2,2,26,...,6,2,2,2. Let $f$ denote the fold count and let $F$ be the number of faces. Our goal is to show that $\eta({\cal C\/})=f^{2}/F\geq 3$. Note that $\cal C$ partitions $\cal T$ into a finite number of monochrome regions with connected interior.
@@
[…]